The number of atoms in 22g of strontium is z . What is the number of bromine molecules in 20g of bromine gas ?
1
Expert's answer
2015-11-24T14:48:55-0500
The atomic mass of strontium is equal 88.
The number of moths of atoms of strontium in 22 g is equal 22/88 = 0,25 mol
On a statement of the problem the mass of 1 atom of strontium is equal 22/Z g
The molecular mass of bromine is equal 160 g.
The number of moths of bromine in 20 g is equal 20/160 = 0,125.
As 1 mol of any substance contains identical quantity of molecules, in 20 g of bromine Z*0.125/0.25 = Z/2 molecules contains.
Answer: Z/2 of molecules